How can i work out a problem by finding the perimeter to the nearest tenth with the area of 90ft2

Found 2 solutions by Anthea Lawn, Theo:
Answer by Anthea Lawn(22)

Depends what shape it is

Answer by Theo(13342)

area of a square? a rectangle? a circle? a parallelogram? a rhombus? a pentagon?
it makes a difference.
in each of them the formula for the area and the formula for the perimeter could be different.
since the only parameter you provided is the area, then assumptions would need to be made as to what was the length, or what was the width, or what was the radius.
you need to be more specific as to what you are referring to.
i'll solve it for a rectangle, so you can see what i mean.
the area of a rectangle is L * W = Area
the perimeter of a rectangle is 2L + 2W = Perimeter
since the area is 90 square feet, then the area formula becomes:
L * W = 90
you don't know what L is and you don't know what W is, so it's difficult to find the perimeter.
this can be solved, but only in terms of one of the dimensions.
since L * W = 90, we can derive:
L = 90/W
we now go to the perimeter formula which states:
2L + 2W = Perimeter
we don't know what the perimeter is, but we can solve for the perimeter if we know either L or W.
we can use L = 90/W to substitute for L in this equation to get:
2L + 2W = P becomes:
2*(90/W) + 2W = P
expand this to get:
180/W + 2W = P
if we know W, we can solve for L and therefore solve for A and for P.
assume W = 5
A = L*W becomes 90 = L*5 which becomes W = 18 because 5*18 = 90
we take these dimensions to the perimeter equation and get:
2L + 2W = P which becomes 2*18 + 2*5 = 36 + 10 = 46
Now suppose W = 10.
W could be anything because we were not told what it should be.
When W = 10, the area equation becomes:
L*10 = 90 which results in L = 9
when L = 9 and W = 10, the perimeter equation becomes:
2*9 + 2*10 = 18 + 20 = 38
we get a different perimeter.
it makes a difference what W is.
we'll get the same area because we use that formula to solve for L.
we don't, however, get the same perimeter.
bottom line is you need to provide more information.
you need to provide what the object is that you have the area of.
you also need to provide at least enough of the dimensions so that the other dimensions can be solved for.